Highlander Cutlery manufactures kitchen knives. One of the employees, whose job is to cut out wooden knife handles, worked 53 hours during a week in January. The employee earns $16 per hour for a 40-hour week, and overtime is paid after 40 hours. For additional hours the employee is paid an overtime rate of $21 per hour. The employee’s time was spent as follows:

Regular duties involving cutting out knife handles 45 hours
General shop cleanup duties 5 hours
Idle time due to power outage 3 hours

Required:

1. Calculate the total cost of the employee’s wages during the week described above.

2. Determine the portion of this cost to be classified in each of the following categories:

a. direct labor

b. manufactoring overhead (idle time)

c. manufacturing overhead (overtime premium)

d manufactoring overhead (indirect labor)

total cost of wages:
Answer #1
  1. Total cost of employees wages will be as follows:
  • Total no. of hours worked is 53 hours.
  • Wages rate for 40 hours is $16 / hour And $ 21 for over time
  • So, total cost will be. (40*16 ) + ( 13*21 ) = (640+273 ) = $ 913.
  1. Portion of cost to be classified in each category is
  • Manufacturing overhead (idle time) = (Total cost/total hours ) X (idle hours ) = ($913/53) X 3 = $ 51.68.
  • Manufacturing overhead (overtime premium ) is numbers of hours exceeding 40 hours multipled by $21.

So, overtime premium = (53-40) X 21 = $273.

  • Manufacturing overhead (indirect wages) = 5 hours of general clean up duties .

So, it will also be calculated similar to idle time cost i.e

(Total cost / total hours) X cleanup hours. = (913/53) X 5 = $86.13.